Calcium hydroxide and calcium sulfate are both used to prepare a saturated solution, S. Both
solids are only sparingly soluble, and the pH of S is found to be 11.6.
(i) Given that the Ksp for calcium hydroxide is 4.0 x 10-5
and the Ksp for calcium sulfate is
2.5 x 10-5
, determine the concentration of sulfate ions present in S.
1
Expert's answer
2020-06-08T15:50:56-0400

"pH=11.6 
\\
pOH =14-11.6=2.4"

let's calculate the concentration of OH- from the pOH

"[OH^-]=4\times10^{-3}"


now calculating the concentration of calcium ions from the Ksp of Calcium hydroxide

"[Ca^{2+}]=\frac{K_{SP}}{[OH^-]^2}=\frac{4\times10^{-5}}{16\times10^{-6}}=2.5"


similarly calculating the concentration of sulfate ions from Ksp of Calcium Sulfate

"[SO_4^{2-}]=\frac{K_{SP}}{Ca^{2+}}=\frac{2.5\times10^{-5}}{2.5}=10^{-5}"